find the area of a rectangle with base of 2yd and height 5ft
Answer by drj(1380)

Find the area of a rectangle with base of 2yd and height 5ft

Step 1. Area A=b*h where b is the base and h is the height

Step 2. 1 yard = 3 feet or 2 yards = 6 feet.

Step 3. A=6*5=30 square feet

Step 4. ANSWER: The area is 30 square feet.

I hope the above steps and explanation were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J
http://www.FreedomUniversity.TV